Assume that the linear cost and revenue models apply. An item costs $4 to make. If fixed costs are $2700 and profits are $13300 when 1000 items are made and sold, find the revenue equation. (Let x be the number of items.)
Answer by ankor@dixie-net.com(22740)

Assume that the linear cost and revenue models apply.
An item costs $4 to make.
If fixed costs are $2700 and profits are $13300 when 1000 items are made and sold, find the revenue equation. (Let x be the number of items.)
:
Cost equation
C(x) = 4x + 2700
:
Rev - cost = 13300 (profit)
let p = the selling price
x = 1000
1000p - (4(1000) + 2700) = 13300
1000p - 4000 - 2700 = 13300
1000p - 6700 = 13300
1000p = 13300 + 6700
1000p = 20000
p = 20000/1000
p = $20 is the selling price
therefore the revenue equation
R(x) = 20x
: